9b) A box contains 10 screws, 3 of which are defective. Two screws are drawn at

random. Find the probability that none of them is defective, if the sample is drawn

(i) with replacement,

(ii) without replacement.


1
Expert's answer
2021-05-10T13:12:13-0400

Total number of screws "= 10"

Defective screws "= 3"

We have to find the probability that none of them is defective

a.) Number of ways of selecting 2 screws from 10 screws with replacement "= ^{10}C_1 \times ^{10}C_1"


Hence, The probability that none of them is defective, if the sample is drawn with replacement


"= \dfrac{^7C_1}{^{10}C_1} .\dfrac{^7C_1}{^{10}C_1}"


"= \dfrac{7}{10}.\dfrac{7}{10}"


"= \dfrac{49}{100}"


b.) Number of ways of selecting 2 screws from 10 screws without replacement "= ^{10}C_2"


Hence, The probability that none of them is defective, if the sample is drawn with replacement



"= \dfrac{^{7}C_2}{^{10}C_2}"


"= \dfrac{42}{90}"


"= \dfrac{21}{45}"